The sum of the digit of a 2-digit number is 7. If 27 is added to the number, the result is expressed by the digits in reverse order. Find the number
Answer by ankor@dixie-net.com(12701)

The sum of the digit of a 2-digit number is 7. If 27 is added to the number,
the result is expressed by the digits in reverse order. Find the number
:
Let x = the 10's digit
Let y = the units digit
:
x + y = 7
y = (7 - x)
:
"If 27 is added to the number, the result is expressed by the digits in reverse order."
10x + y + 27 = 10y + x
10x - x = 10y - y
9x + 27 = 9y
Simplify divide by 9
x + 3 = y
Substitute for y
x + 3 = 7 - x
x + x = 7 - 3
2x = 4
x = 2
then
y = 7-2
y = 5
:
the number: 25
:
Check solution in the statement:
"If 27 is added to the number, the result is expressed by the digits in reverse order."
25 + 27 = 52